# Efficient Hamiltonian, structure and trace distance learning of Gaussian states

### v2

@@ 6 Sparse approximations of $(2V-i\Omega)^{-1}$
$\|M\|_{\infty}\|M^{-1}\|_{\infty}$ is the condition number of $M$. Looking at the proof, it is clear that in instead of $\|M\|_{\infty}\|M^{-1}\|_{\infty}$ we can leave $\|D^{-1}C\|_{\infty}$ in the bound statement. When $C$ contains very few columns with respect to the size of $D$, and $D$ has a large condition number, it would be natural to guess that $\|D^{-1}C\|_{\infty}$ does not scale as the norm of $\|D^{-1}\|$, as the subspace spanned by the column space of $C$ may be not aligned with the eigenspaces with largest eigenvalues of $\|D^{-1}\|$. All the following bounds can be in principle improved using this fact. We prefer not to do it explicitly since a bound on $\|D^{-1}C\|_{\infty}$ is not a natural assumption in our context.
@@
$\leq\max_{i\in[m]}4\|M\|_{\infty}^{2}|\mathcal{N}_{i}|\zeta\,,$
@@ 8 Learning the graph
$\zeta$ $=\min\left\{\frac{\epsilon^{\prime}}{C^{\prime}(s,\beta_{\max})\Delta^{2l}},\,\frac{\epsilon^{\prime}}{e^{2\beta_{\max}}18\,\Delta^{2l}s^{4}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{(1-e^{-2\beta_{\min}})}{(1-e^{-2\beta_{\max}})}},\,\frac{e^{-2\beta_{\max}}}{2\Delta^{l}s^{2}}\right\}$
@@
[…]
  • When ${\mathcal{N}}_{i}$ coincides with the $l$-neighborhood $\mathcal{N}(l)_{i}$ (see Definition 6.5), we assume that $\max_{\delta\in\{0,1\}}\|B_{i,\delta}-\widetilde{B}_{i,\delta}\|_{\infty}\|M^{-1}\|_{\infty}\|M\|_{\infty}\leq\eta\,,$ and at the same time $\|E\|_{\infty}\leq\eta$ for the error $E$ defined in Lemma 6.1, which we can always ensure by taking $l$ large enough. In fact, it suffices to impose that
@@
after some simplifications, where we used Stirling’s lower bound $n!\geq n^{n}/e^{n}$ and denoted $C(S,d_{\min},d_{\max}):=18\|S\|_{\infty}^{10}e^{6d_{\max}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\frac{1-e^{-2d_{\max}}}{1-e^{-2d_{\min}}}\right)^{2}}$. Then, $2\xi(l)^{2}\zeta\|S\|^{4}_{\infty}e^{4d_{\max}}\leq\epsilon^{\prime}$, implying that $\|M-\operatorname{LI}_{\hat{\mathcal{N}}}(2\hat{V}-i\Omega)\|_{\infty}\leq 2\epsilon^{\prime}$, $\zeta\leq 2\eta$ and $\zeta\leq\frac{1}{4\xi(l)}\|(2V-i\Omega)^{-1}\|_{\infty}^{-1}$, are ensured by taking
$\zeta$ $=\min\left\{\frac{\epsilon^{\prime}}{C^{\prime}(S,d_{\max})\Delta^{2l}},\,\frac{\epsilon^{\prime}}{e^{2d_{\max}}18\,\Delta^{2l}\|S\|_{\infty}^{4}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1-e^{-2d_{\min}}}{1-e^{-2d_{\max}}}},\,\frac{e^{-2d_{\max}}}{2\Delta^{l}\|S\|_{\infty}^{2}}\right\}\,,$
with $C^{\prime}(S,d_{\max}):=2\|S\|_{\infty}^{4}e^{4d_{\max}}$, where the last bound is implied by (70). Now, we want $\epsilon^{\prime}$ to be small enough so that $\|\hat{H}-H\|_{\infty}\leq\kappa/2$, where we compute $\hat{H}$ as in Equation 37 from $\operatorname{LI}_{\hat{\mathcal{N}}}(2\hat{V}-i\Omega)$. As in Remark 6.4 we can take $\epsilon^{\prime}=\min\left(\kappa\frac{e^{-2d_{\max}}\|S\|^{-2}_{\infty}}{16\left(1+{e^{4d_{\max}}\|S\|_{\infty}^{4}(1-e^{-2d_{\max}})}\right)},\frac{e^{-2d_{\max}}\|S\|_{\infty}^{-2}}{8}\right)$. Since now we ensured that $\|\hat{H}-H\|_{\infty}\leq\kappa/2$, we can learn the correct edges. By Lemma 4.1, the number of copies we used to get $\hat{V}$ satisfying $|\hat{V}_{k,k^{\prime}}-V_{k,k^{\prime}}|\leq\zeta/2$ is

### v3

@@ 6 Sparse approximations of $(2V-i\Omega)^{-1}$
$\|M\|_{\infty}\|M^{-1}\|_{\infty}$ is the condition number of $M$. Looking at the proof, it is clear that instead of $\|M\|_{\infty}\|M^{-1}\|_{\infty}$ we can leave $\|D^{-1}C\|_{\infty}$ in the bound statement. When $C$ contains very few columns with respect to the size of $D$, and $D$ has a large condition number, it would be natural to guess that $\|D^{-1}C\|_{\infty}$ does not scale as the norm of $\|D^{-1}\|$, as the subspace spanned by the column space of $C$ may be not aligned with the eigenspaces with largest eigenvalues of $\|D^{-1}\|$. All the following bounds can be in principle improved using this fact. We prefer not to do it explicitly since a bound on $\|D^{-1}C\|_{\infty}$ is not a natural assumption in our context.
@@
$\leq\max_{i\in[m]}4\|M\|_{\infty}^{2}|\mathcal{N}_{i}|\zeta\,,$ (36)
@@ 8 Learning the graph
$\zeta$ $=\min\left\{\frac{\epsilon^{\prime}}{C^{\prime}(s,\beta_{\max})\Delta^{2l}},\,{\frac{\epsilon^{\prime}e^{-4\beta_{\max}}}{72\,\Delta^{3l}s^{6}}\frac{1-e^{-2\beta_{\min}}}{1-e^{-2\beta_{\max}}}}\right\}$
@@
  • For any $j\in{\widetilde{\mathcal{N}}}_{i}$, $\delta_{1},\delta_{2}\in\{0,1\}$, we have $|\operatorname{LI}_{\widetilde{\mathcal{N}}}(2V-i\Omega)_{2i-\delta_{1},2j-\delta_{2}}-\operatorname{LI}_{\widetilde{\mathcal{N}}}(2\hat{V}-i\Omega)_{2i-\delta_{1},2j-\delta_{2}}|\leq\eta$. We can ensure this by increasing the number of copies of $\rho(H,t)$ so that each entry of $\hat{V}$ is close to the corresponding entry of $V$ with high probability. By the argument of (36), with $N=2V-i\Omega$, $\hat{N}=2\hat{V}-i\Omega$ and $M=(2V-i\Omega)^{-1}$, if $|2\hat{V}_{k,k^{\prime}}-2V_{k,k^{\prime}}|\leq{\zeta}$, with $\zeta\leq\frac{1}{8\xi(l)}\|(2V-i\Omega)^{-1}\|_{\infty}^{-1}$ we get $|\operatorname{LI}_{\widetilde{\mathcal{N}}}(2V-i\Omega)_{2i-\delta_{1},2j-\delta_{2}}-\operatorname{LI}_{\widetilde{\mathcal{N}}}(2\hat{V}-i\Omega)_{2i-\delta_{1},2j-\delta_{2}}|\leq 8\xi(l)\|(2V-i\Omega)^{-1}\|_{\infty}\zeta$. Thus, we need to ensure also $\zeta\leq\frac{\|(2V-i\Omega)^{-1}\|_{\infty}^{-1}\eta}{8\xi(l)}$.
  • When ${\mathcal{N}}_{i}$ coincides with the $l$-neighborhood $\mathcal{N}(l)_{i}$ (see Definition 6.5), we assume that
    $\max_{\delta\in\{0,1\}}\|B_{i,\delta}-\widetilde{B}_{i,\delta}\|_{\infty}\|M^{-1}\|_{\infty}\|M\|_{\infty}\leq\eta\,,$ and at the same time $\|E\|_{\infty}\leq\eta$ for the error $E$ defined in Lemma 6.1, which we can always ensure by taking $l$ large enough. In fact, it suffices to impose that
@@
after some simplifications, where we used Stirling’s lower bound $n!\geq n^{n}/e^{n}$ and denoted $C(S,d_{\min},d_{\max}):=18\|S\|_{\infty}^{10}e^{6d_{\max}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left(\frac{1-e^{-2d_{\max}}}{1-e^{-2d_{\min}}}\right)^{2}}$. Then, $2\xi(l)^{2}\zeta\|S\|^{4}_{\infty}e^{4d_{\max}}\leq\epsilon^{\prime}$ implies that $\|M-\operatorname{LI}_{\hat{\mathcal{N}}}(2\hat{V}-i\Omega)\|_{\infty}\leq 2\epsilon^{\prime}$. Moreover, assuming $\eta\leq 1$, $\zeta\leq 2\eta$, $\zeta\leq\frac{1}{8\xi(l)}\|(2V-i\Omega)^{-1}\|_{\infty}^{-1}\leq\frac{1}{4\xi(l)}\|(2V-i\Omega)^{-1}\|_{\infty}^{-1}$, and ${\zeta\leq\frac{\|(2V-i\Omega)^{-1}\|_{\infty}^{-1}\eta}{8\xi(l)}}$ are ensured by simply the last inequality, which is implied by taking ${\zeta\leq\frac{e^{-2d_{\max}}\eta}{4\|S\|_{\infty}^{2}\xi(l)}}$, as a consequence of (71) (this also makes the assumption $\eta\leq 1$ verified). One can thus choose
$\zeta$ $=\min\left\{\frac{\epsilon^{\prime}}{C^{\prime}(S,d_{\max})\Delta^{2l}},\,{\frac{\epsilon^{\prime}e^{-4d_{\max}}}{72\,\Delta^{3l}\|S\|_{\infty}^{6}}\frac{1-e^{-2d_{\min}}}{1-e^{-2d_{\max}}}}\right\}\,,$
with $C^{\prime}(S,d_{\max}):=2\|S\|_{\infty}^{4}e^{4d_{\max}}$. Now, we want $\epsilon^{\prime}$ to be small enough so that $\|\hat{H}-H\|_{\infty}\leq\kappa/2$, where we compute $\hat{H}$ as in Equation 38 from $\operatorname{LI}_{\hat{\mathcal{N}}}(2\hat{V}-i\Omega)$. As in Remark 6.4 we can take $\epsilon^{\prime}=\min\left(\kappa\frac{e^{-2d_{\max}}\|S\|^{-2}_{\infty}}{16\left(1+{e^{4d_{\max}}\|S\|_{\infty}^{4}(1-e^{-2d_{\max}})}\right)},\frac{e^{-2d_{\max}}\|S\|_{\infty}^{-2}}{8}\right)$. Since now we ensured that $\|\hat{H}-H\|_{\infty}\leq\kappa/2$, we can learn the correct edges. By Lemma 4.1, the number of copies we used to get $\hat{V}$ satisfying $|\hat{V}_{k,k^{\prime}}-V_{k,k^{\prime}}|\leq\zeta/2$ is